OpenStudy (thatonegirl_):

How would you integrate 1+lnx ?

zepdrix (zepdrix):

lnx isn't super obvious just by looking at it. It will require integration by parts.

zepdrix (zepdrix):

Get the 1 out of the way so we can ignore it.\[\large\rm \int\limits 1+\ln x~dx\quad=\quad x+\int\limits \ln x~dx\]

zepdrix (zepdrix):

\[\large\rm u=\ln x\qquad\qquad\qquad dv=dx\]Integrating lnx would be... hmm weird. That's what we're trying to get around anyway. So we want to differentiate our log instead. Hence, it is our u.

OpenStudy (thatonegirl_):

What's dv?

zepdrix (zepdrix):

It's the invisible 1 next to the ln(x). ln(x) = 1*ln(x) We're integrating the 1.

OpenStudy (thatonegirl_):

oh ok

zepdrix (zepdrix):

It's better to get comfortable with differentials though. dv = dx If that's confusing though, and you'd rather look at the function stuff, then yes, it's the 1.

zepdrix (zepdrix):

\[\large\rm u=\ln x\qquad\qquad\qquad dv=dx\]So we get our other pieces that we need,\[\large\rm du=\frac1x~dx\qquad\qquad\qquad v=x\]

OpenStudy (thatonegirl_):

ok looks good

zepdrix (zepdrix):

So umm.. one of things you might notice is... the log turns into a power of x when it's differentiated. That will mix nicely with our new v showing up.

zepdrix (zepdrix):

So we get something like this from our parts, yes?\[\large\rm \int\limits \ln x~dx\quad=\quad x \ln x-\int\limits x\cdot\frac{1}{x}~dx\]Do you see how that new integral that shows up is much easier to deal with now?

OpenStudy (thatonegirl_):

ohh ok i see

zepdrix (zepdrix):

\[\large\rm \int\limits\limits \ln x~dx\quad=\quad x \ln x-\int\limits dx\]

zepdrix (zepdrix):

\[\large\rm \int\limits\limits\limits \ln x~dx\quad=\quad x \ln x-x+c\]

zepdrix (zepdrix):

And then don't forget about the +1 that we integrated earlier on, right?

OpenStudy (thatonegirl_):

oh yeah so that would then cancel out this x right?

zepdrix (zepdrix):

\[\large\rm \int\limits(1+\ln x)~dx\quad=\quad x \ln x+c\]Yayyy good job \c:/

OpenStudy (thatonegirl_):

thank you!!! This helped a lot!!!!

zepdrix (zepdrix):

yay team